Suppose today is February 10, 2017, and your firm produces breakfast cereal and needs 170,000 bushels of corn in May 2017 forTABLE 23.1 Sample Wall Street Journal Futures Price Quotations Chg Open interest -0.0125 -0.0130 6 72 120,434 -2.50 -2.70 -2.-1.25 -1.25 500,405 346,739 50 3.00 4,155 3,027 Agriculture Futures Corn (CBT)-5,000 bu; cents per bu. March 369.50 372.75 36-1-030 -1-030 622,929 2,310 -14.0 -14.5 3,169,794 68,648 Interest Rate Futures Treasury Bonds (CBT)-$100,000; pts 32nds of 10126,519 893 68,270 1,259 2,882,114 36,548 Index Futures Mini DJ Industrial Average (CBT)-$5 x index March 20010 2 0158 A 1998
Answer #1

1.
Total cost=170000*377.25/100
=641325.00

2.
Loss of 32725.00

Profit=170000*3.58-170000*377.25/100=-32725.0000